OpenStudy (anonymous):

Greg has two porcelain spheres. The radii of the spheres are in a ratio of 3 : 4. What is the ratio between the volumes of the two spheres? A.) 9 : 16 B.) 27 : 64 C.) 9 : 64 D.) 27 : 32

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

2 spheres -> similar figures For similar figures, (ratio of sides)^3 = ratio of volumes So, \[(\frac{3}{4})^3 = ratio\ of\ volume\] Can you simplify it and get the answer?

OpenStudy (anonymous):

27/64?

OpenStudy (callisto):

Yes~

OpenStudy (callisto):

It should be :)

OpenStudy (anonymous):

ok thank you :)

OpenStudy (callisto):

Welcome :)